Question 5 2 pts Lion Inc., has sales of $2550, total assets of $1021, and a debt-equity ratio of 0.7. if its return on equity is 13%. What is Lions Net Income? (Round final answer to 2 decimal places. Do not round intermediate calculations). Topic: Financial Ratios (DuPont Model)
Answer #1
The Dupont Model equates ROE to profit margin, asset turnover, and financial leverage.
The formula to calculate ROE i.e.return on equity = Profit margin ratio x Total asset turnover ratio x Financial leverage ratio
Let us assume Net Income be a.
Profit margin ratio = Net Income / Sales = a / $2550
Total asset turnover ratio = Sales / Total assets = $2550 / $1021 = 2.50
Total assets = Debt + equity = 0.70 + 1 = 1.70
Financial leverage ratio = Total assets / Total equity = 1.70/1 = 1.70
Putting above value in a dupont formula
0.13 = [a/$2550] x 2.50 x 1.70
0.03062 = [a/$2550]
a = 78.08
Lion's net income = $78.08